PROBLEM 2. The following information was taken from the books and records of Legg, Inc.:

- Net Income   =   $480,000

- Capital structure:

A. Convertible 6% bonds. Each of the 300, $1,000 bonds is convertible

into 50 shares of common stock at the present date and for the next

10 years. =   $300,000

   B. $10 par common stock, 200,000 shares issued and outstanding

during the entire year. =    $2,000,000

  C. Stock warrants outstanding to buy 16,000 shares of common stock

at $20 per share.

- Other information:

A. Bonds converted during the year =  NONE

B. Income tax rate =  30%

C. Convertible debt was outstanding the entire year

D. Average market price per share of common stock during the year =  $32

   E. Warrants were outstanding the entire year

   F.    Warrants exercised during the year =  NONE

Instructions

1. Compute basic and diluted earnings per share.
Answer #1

Earnings per share is calculated by dividing the earnings available for distribution by total number of weighted average shares.

On the other hand, Diluted EPS takes into potential issue of shares. Each potential share issue has its own merits and viewed accordingly on individual basis.

Basic EPS = Earnings available to shareholders/Number of shares

                   = 480,000/200.000 = $ 2.40 per share

Diluted EPS

Security

Net income

Adjustment

Adjusted
net income

shares

Adjustment

Adjusted
shares

Diluted EPS

Com stock

480,000

480,000

200,000

200,000

2.4

Warrants

480,000

480,000

200,000

6000*

206000

2.33

Conv bonds

480,000

12600*

492,600

200,000

15000*

221000

2.23

Number of shares at average market price = 320,000/32 = 10,000

Stock warrants outstanding = 16000

Adjustment needed = 6,000

Convertible bonds

Adjustment in net income = 300,000 * 6% * (1-tax) = $ 12600

Adjustment needed in shares = 300 * 50 = 15000 shares